OpenStudy (anonymous):

Halley's comet has an elliptical orbit with the sun at one focus. its orbit is given below approx. by r=10.21/1+0.874sin theta. in the formula r is measure in astronomical units. (one astronomical unit is the average distance from the Earth to the sun, approx. 93 million miles.) Whats the distance from Halley's comet to the sun at the shortest distance from the sun?

OpenStudy (wolf1728):

I'm an amateur astronomer (one of the zillions) and I'm somewhat familiar with the mathematics of orbits but I'll admit I've never seen an orbital formula such as: r=10.21/1+0.874sin theta Is that just one formula? Is it exactly the way the problem is stated? Don't think I'm just giving you a hard time about this. I really want to help.

OpenStudy (anonymous):

Halley's comet has an elliptical orbit with the sun at one focus. its orbit is given approximately by r=10.21/1+0.874sin theta. in the formula r is measure in astronomical units. (one astronomical unit is the average distance from the Earth to the sun, approx. 93 million miles.) Find the distance from Halley's comet to the sun at the shortest distance from the sun?

OpenStudy (anonymous):

i changed two words

OpenStudy (anonymous):

thank you

OpenStudy (wolf1728):

How exactly am I supposed to read that formula? r=10.21/1+0.874sin theta Among other things it mentions the sine of theta. Shouldn't we be told what theta is? 10 degrees? 2 radians?

OpenStudy (ranga):

Is the formula \(\Large r = \frac{10.21}{1+0.874sin(\theta)} \)?

OpenStudy (anonymous):

yes

OpenStudy (ranga):

Then the above formula is probably in polar coordinates with r measured from the focus. r will be minimum when \(sin(\theta)\) is maximum which is 1.

OpenStudy (anonymous):

I think answer is 11.68 astronomical units, 1086 million miles

OpenStudy (anonymous):

im not sure though

OpenStudy (anonymous):

i've been studying polar coordinates so that sound right Ranga

OpenStudy (wolf1728):

no the perihelion is .586 astronomical units.

OpenStudy (anonymous):

How many millions of miles?

OpenStudy (anonymous):

543 million Miles?

OpenStudy (wolf1728):

Wkipedia says aphelion = 35.1 AU perihelion = .586 AU My calculations: average = 17.843 AU eccentricity = (aph -per) / (aph + per) = 0.9671579891 1 au is about 92,900,000 miles so .586 au = 54,439,400 miles

OpenStudy (anonymous):

thank you wolf

OpenStudy (wolf1728):

it's okay u r welcome :-)